# Morita equivalence of Nijenhuis structures

### v1

@@ 3.3 Morita equivalence of infinitesimal structures
We show that this condition is equivalent to Equations (35) and (36) by using the identification (39) between related $(1,1)$-tensors and related $1$-derivations. Let $\left(\mu^{*}\nabla,\mu^{*}\ell,J\right)$ be the $1$-derivation associated with $\mu^{*}(R_{1}\times R_{2})$ and recall that the $1$-derivation corresponding to $J^{\mathrm{tg}}$ is $(\nabla^{J},J,J)$. Due to (39), the condition that $\mu^{*}(R_{1}\times R_{2})$ is $\mathscr{a}$-related to $J^{\mathrm{tg}}$ is equivalent to:
@@
3.4 Lie correspondence for Morita equivalences
@@
Consider the graph of $\mathscr{A}$ as a Lie groupoid $\mathscr{K}=\left(\mathrm{Graph}(\mathscr{A})\rightrightarrows P\right)$, where $\mathrm{Graph}(\mathscr{A})\subseteq P\times_{(\mu,\mathbf{t})}(G_{1}\times H_{1})\times_{(\mu,\mathbf{s})}P$. The source and target maps are the respective projections onto the second and first $P$-factors; thus, we may view $\mathscr{K}$ as a subgroupoid of $\mathscr{G}\times\mathscr{H}\times\mathrm{Pair}(P)$. The condition that the tensor fields $J$ and $(N_{1}\times N_{2})\times J$ intertwine the action (Equation (7)) implies that $N\coloneqq J\times(N_{1}\times N_{2})\times J$ is a morphism from the tangent Lie groupoid $T\mathscr{K}$ to itself. Thus, $N\in\Omega^{1}(\mathscr{K},T\mathscr{K})$ is a multiplicative $(1,1)$-tensor field on $\mathscr{K}$ covering $J$.
The Lie algebroid of $\mathscr{K}=\left(\mathrm{Graph}(\mathscr{A})\rightrightarrows P\right)$ is the graph of the infinitesimal action $\mathscr{a}:\Gamma\left(A_{\mathscr{G}}\times A_{\mathscr{H}}^{\mathrm{op}}\right)\to\mathfrak{X}(P)$; that is, $A_{\mathscr{K}}=\mathrm{Graph}(\mathscr{a})$. The Lie algebroid $A_{\mathscr{K}}$ is a Lie subalgebroid of $(A_{\mathscr{G}}\times A_{\mathscr{H}}^{\mathrm{op}})\times_{(\rho,\mathrm{d}\mu)}TP\to P$, which integrates to $\mathscr{G}\times\mathscr{H}^{\mathrm{op}}\times\mathrm{Pair}(P)$. Applying the Lie functor to the morphism $N=J\times(N_{1}\times N_{2})\times J$ and using Lemma 3.6 to identify $J^{\mathrm{tg}}$ as the infinitesimal counterpart of the factor $J\times J$, we obtain a morphism of Lie algebroids $N_{*}=N_{1*}\times N_{2*}\times J^{\mathrm{tg}}:TA_{\mathscr{K}}\to TA_{\mathscr{K}}$. The condition that $N_{*}$ preserves $\mathrm{Graph}(\mathscr{a})$ is equivalent to
$\mathrm{d}\mathscr{a}\circ\mu^{*}\left(N_{1*}\times N_{2*}\right)=J^{\mathrm{tg}}\circ\mathrm{d}\mathscr{a}.$
This equation is one of the requirements of Definition 3.12 (cf. Remark 3.13). It remains to observe that the action $\mathscr{a}$ decomposes as the sum $\mathscr{a}=\mathscr{a}_{1}+\mathscr{a}_{2}$, where $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ are the differentiations of $\mathscr{A}_{1}$ and $\mathscr{A}_{2}$, respectively. By Lemma 3.10, $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ constitute an infinitesimal Morita $\left(A_{\mathscr{G}},A_{\mathscr{H}}\right)$-bibundle $P$. This completes the proof for multiplicative $(1,1)$-tensor fields.
Regarding the Nijenhuis property, the vanishing of the torsion for the multiplicative tensor fields implies that the IM tensor fields are also Nijenhuis (see [9, Section 6.2]). ∎
@@
Let $\mathscr{G}$ and $\mathscr{H}$ be source-simply connected Lie groupoids with Lie algebroids $A_{\mathscr{G}}$ and $A_{\mathscr{H}}$ respectively, and let $P$ be an infinitesimal Morita $(A_{\mathscr{G}},A_{\mathscr{H}})$-bibundle. Let $R_{1}\in\Omega^{1}(A_{\mathscr{G}},TA_{\mathscr{G}})$ and $R_{2}\in\Omega^{1}(A_{\mathscr{H}},TA_{\mathscr{H}})$ be IM tensor fields that are Morita equivalent via $(P,J)$ for some $J\in\Omega^{1}(P,TP)$. Then the multiplicative $(1,1)$-tensor fields $R_{\mathscr{G}}$ and $R_{\mathscr{H}}$ that integrate $R_{1}$ and $R_{2}$ are Morita equivalent via the associated Morita $(\mathscr{G},\mathscr{H})$-bibundle $(P,J)$. Furthermore, if the infinitesimal Morita equivalence is Nijenhuis, then the global Morita equivalence is also Nijenhuis.
@@
We denote the respective left and right infinitesimal actions of the Lie algebroids by
$\mathscr{a}_{1}:\mu_{1}^{*}A_{\mathscr{G}}\to TP\qquad\text{and}\qquad\mathscr{a}_{2}:\mu_{2}^{*}A_{\mathscr{H}}\to TP.$
The sum of the actions is then $\mathscr{a}(\xi,\zeta)=\mathscr{a}_{1}(\xi)+\mathscr{a}_{2}(\zeta)$. By Remark 3.13, the condition for the tensor fields to be Morita equivalent is that the image of $T\mathrm{Graph}(\mathscr{a})$ under the map $R\coloneqq\mu^{*}(R_{1}\times R_{2})\times J^{\mathrm{tg}}$ is contained in $T\mathrm{Graph}\left(\mathscr{a}\right)$, where $\mu\coloneqq(\mu_{1},\mu_{2})$. Moreover, since each factor of $R$ is a Lie algebroid morphism, $R$ defines a Lie algebroid morphism from $T\mathrm{Graph}\left(\mathscr{a}\right)$ to itself.
Since the Lie algebroids are integrable and the infinitesimal actions $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ are complete, we integrate them to global actions of the Lie groupoids, $\mathscr{A}_{1}:G_{1}\times_{G_{0}}P\to P$ and $\mathscr{A}_{2}:P\times_{H_{0}}H_{1}\to P$. These global actions realize a Morita bibundle $G_{0}\xleftarrow{\ \mu_{1}\ }P\xrightarrow{\ \mu_{2}\ }H_{0}$ between the Lie groupoids $\mathscr{G}$ and $\mathscr{H}$.
Let $\mathscr{A}:(G_{1}\times H_{1})\times_{(\mathbf{s},\mu)}P\to P$ denote the combined action $\mathscr{A}\left(g,h,p\right)=g\cdot p\cdot h$, which integrates $\mathscr{a}$. The graph of this action is a Lie subgroupoid $\mathscr{K}\coloneqq\left(\mathrm{Graph}(\mathscr{A})\rightrightarrows P\right)$ of $\mathscr{G}\times\mathscr{H}^{\mathrm{op}}\times\mathrm{Pair}(P)$.
Observe that $\mathscr{K}$ is a source-simply connected Lie groupoid since its source-fibers are products of simply-connected components. Consequently, $T\mathscr{K}$ is source-simply connected.
Using Lie’s second theorem, we integrate the Lie algebroid morphism $R:T\mathrm{Graph}(\mathscr{a})\to T\mathrm{Graph}(\mathscr{a})$ to a Lie groupoid morphism $R_{\mathscr{K}}:T\mathscr{K}\to T\mathscr{K}$. Thus, $R_{\mathscr{K}}$ is a multiplicative $(1,1)$-tensor field on $\mathscr{K}$. Let $R_{\mathscr{G}}\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ and $R_{\mathscr{H}}\in\Omega^{1}(\mathscr{H},T\mathscr{H})$ be the integrations of each of the IM $(1,1)$-tensor fields $R_{1}$ and $R_{2}$. Since $J^{\text{tg}}$ integrates to $J\times J$ (Lemma 3.6), the uniqueness of integration implies that $R_{\mathscr{K}}$ coincides with $J\times(R_{\mathscr{G}}\times R_{\mathscr{H}})\times J$ restricted to the graph of $\mathscr{A}$.
The condition that $J\times(R_{\mathscr{G}}\times R_{\mathscr{H}})\times J$ preserves $\mathrm{Graph}(\mathscr{A})$ is equivalent to the statement that $G_{0}\xleftarrow{\ \mu_{1}\ }\left(P,J\right)\xrightarrow{\ \mu_{2}\ }H_{0}$ is a Morita equivalence between the multiplicative tensor fields $R_{\mathscr{G}}$ and $R_{\mathscr{H}}$. Since the vanishing of the Nijenhuis torsion is preserved through integration (see [9, Section 6.2]), this completes the proof. ∎
@@
4 Compatibility with quasi-symplectic groupoids and Dirac structures
In this section, we apply the framework of Morita equivalence for multiplicative $(1,1)$-tensor fields, developed in Section 2, to Lie groupoids equipped with geometric structures compatible with the Nijenhuis condition. Building upon the established notion of Morita equivalence for quasi-symplectic groupoids introduced by Xu [41], our primary objective is to extend this definition to quasi-symplectic-Nijenhuis groupoids and prove that it constitutes an equivalence relation (Theorem 4.6).
Subsequently, we investigate the infinitesimal counterpart of this theory as an application of Section 3. Recall that a Morita equivalence of quasi-symplectic groupoids differentiates to a dual pair of Dirac structures satisfying additional conditions [17]. We enhance this correspondence by imposing compatibility with infinitesimal Nijenhuis structures, leading to a notion of Morita equivalence for Dirac-Nijenhuis structures (Definition 4.14). Finally, we establish the consistency of this framework via the Lie functor: Theorems 4.17 and 4.18 prove the global-to-infinitesimal correspondence.
4.1 Quasi-symplectic-Nijenhuis groupoids
We begin by recalling that a quasi-symplectic groupoid [41, 6] (called twisted presymplectic groupoid in [6]) is a triple $(\mathscr{G},\omega,\eta)$, where $\mathscr{G}=\left(G_{1}\rightrightarrows G_{0}\right)$ is a Lie groupoid such that $\operatorname{dim}G_{1}=2\operatorname{dim}G_{0}$, $\omega\in\Omega^{2}(G_{1})$ is a multiplicative $2$-form and $\eta\in\Omega^{3}(G_{0})$ is a closed $3$-form satisfying
$\mathrm{d}\omega=\mathbf{s}^{*}\eta-\mathbf{t}^{*}\eta\qquad\text{and}\qquad\operatorname{ker}\omega\cap\operatorname{ker}\mathrm{d}\mathbf{s}\cap\operatorname{ker}\mathrm{d}\mathbf{t}=0.$ (40)
Although Morita equivalence for quasi-symplectic groupoids can be formulated in terms of Morita spans (see [30, 13]), we give preference to the original definition [41] here. This is because the formulation via spans does not naturally equip the intertwining groupoid with a multiplicative $2$-form, a feature which is essential for stating the compatibility with the Nijenhuis tensor field in our framework.
Recall that a Morita equivalence of quasi-symplectic groupoids [41] $\left(\mathscr{G},\omega_{1},\eta_{1}\right)$ and $\left(\mathscr{H},\omega_{2},\eta_{2}\right)$ is given by a Morita bibundle $G_{0}\leftarrow P\to H_{0}$ equipped with a $2$-form $\varpi\in\Omega^{2}(P)$,
@@ 4.1 Quasi-symplectic-Nijenhuis groupoids
where $\mathscr{A}\colon G_{1}\times_{G_{0}}P\times_{H_{0}}H_{1}\to P$ is the action $(g,p,h)\mapsto g\cdot p\cdot h$. This definition extends the notion of Morita equivalence for symplectic groupoids introduced by Xu [40].
@@
The Morita equivalence of quasi-symplectic groupoids is indeed an equivalence relation [41, Theorem 4.5]. For transitivity, given two principal bibundles $G_{0}\leftarrow(P_{1},\varpi_{1})\to H_{0}$ and $H_{0}\leftarrow(P_{2},\varpi_{2})\to G_{0}^{\prime}$, a $2$-form $\varpi\in\Omega^{2}(P)$ is obtained on the quotient $P=P_{1}\diamond P_{2}=\left(P_{1}\times_{H_{0}}P_{2}\right)/\mathscr{H}$, realizing a Morita equivalence $G_{0}\leftarrow(P,\varpi)\to G_{0}^{\prime}$. The $2$-form $\varpi$ is defined by the property
$\pi^{*}\varpi=\mathrm{pr}_{1}^{*}\varpi_{1}+\mathrm{pr}_{2}^{*}\varpi_{2},$ (41)
@@
To define quasi-symplectic-Nijenhuis groupoids we require a compatibility between $2$-forms and $(1,1)$-tensor fields. Consider the following notation. For any $p$-form $\alpha\in\Omega^{p}\left(M\right)$ and vector-valued $1$-form $K\in\Omega^{1}(M,TM)$, write $\alpha_{K}$ to denote
@@
A $2$-form $\omega\in\Omega^{2}(M)$ and a $(1,1)$-tensor field $K\in\Omega^{1}(M,TM)$ on a manifold $M$ are said to be compatible [28] if
$\omega^{\flat}\circ K=K^{*}\circ\omega^{\flat}\qquad\text{and}$ (42)
$\mathrm{d}(\omega_{K})=(\mathrm{d}\omega)_{K}.$ (43)
@@
A quasi-symplectic-Nijenhuis groupoid is a quadruple $(\mathscr{G},\omega,\eta,N)$, where $(\mathscr{G},\omega,\eta)$ is a quasi-symplectic groupoid and $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ is a multiplicative Nijenhuis tensor field that is compatible with the $2$-form $\omega$.
@@
Following the terminology of [8], a holomorphic presymplectic groupoid is a holomorphic groupoid $(\mathscr{G},I_{\mathscr{G}})$ equipped with a multiplicative holomorphic closed $2$-form $\Omega\in\Omega^{(2,0)}(G_{1})$ that satisfies $\operatorname{ker}\Omega\cap\operatorname{ker}\mathrm{d}\mathbf{s}\cap\operatorname{ker}\mathrm{d}\mathbf{t}=0$. Equivalently, a holomorphic presymplectic groupoid is described as a quasi-symplectic-Nijenhuis groupoid $(\mathscr{G},\omega,0,I_{\mathscr{G}})$, where $\mathrm{d}\omega=0$ and $\Omega=\omega-i\omega_{I_{\mathscr{G}}}$ [8, Section 6.3].
Definition 4.4.
A Morita equivalence of quasi-symplectic-Nijenhuis groupoids $(\mathscr{G},\omega_{1},\eta_{1},N_{1})$ and $(\mathscr{H},\omega_{2},\eta_{2},N_{2})$ is given by a Morita bibundle $G_{0}\leftarrow P\to H_{0}$ equipped with a $2$-form $\varpi\in\Omega^{2}(P)$ and a $(1,1)$-tensor field $J\in\Omega^{1}(P,TP)$,
@@
Two holomorphic presymplectic groupoids $(\mathscr{G},\Omega_{1},I_{\mathscr{G}})$ and $(\mathscr{H},\Omega_{2},I_{\mathscr{H}})$ are Morita equivalent if and only if, there is a holomorphic Morita bibundle $G_{0}\xleftarrow{\ \mu_{1}\ }(P,I_{P})\xrightarrow{\ \mu_{2}\ }H_{0}$ (Example 2.10) and a holomorphic closed $2$-form $\Omega_{P}=\varpi-i\varpi_{I_{P}}\in\Omega^{(2,0)}(P)$ satisfying $\mathscr{A}^{*}\varpi=\mathrm{pr}_{G_{1}}^{*}\omega_{1}+\mathrm{pr}_{P}^{*}\varpi+\mathrm{pr}_{H_{1}}^{*}\omega_{2}$ and $\operatorname{ker}\varpi\cap\operatorname{ker}\mathrm{d}\mu_{1}\cap\operatorname{ker}\mathrm{d}\mu_{2}=0$, where $\Omega_{1}=\omega_{1}-i\omega_{1\,I_{\mathscr{G}}}$ and $\Omega_{2}=\omega_{2}-i\omega_{2\,I_{\mathscr{H}}}$.

### v2

@@ 3.3.1. Infinitesimal Morita equivalence of linear $(1,1)$-tensor fields
We show that this condition is equivalent to Equations (35) and (36) by using the identification (39) between related $(1,1)$-tensors and related 11-derivations. Let $\left(\mu^{*}\nabla,\mu^{*}\ell,J\right)$ be the 11-derivation associated with $\mu^{*}(R_{1}\times R_{2})$ and recall that the 11-derivation corresponding to $J^{\mathrm{tg}}$ is $(\nabla^{J},J,J)$. Due to (39), the condition that $\mu^{*}(R_{1}\times R_{2})$ is $\mathscr{a}$-related to $J^{\mathrm{tg}}$ is equivalent to:
@@
3.4. Lie correspondence for Morita equivalences
@@
Consider the graph of $\mathscr{A}$ as a Lie groupoid $\mathscr{K}=\left(\mathrm{Graph}(\mathscr{A})\rightrightarrows P\right)$, where $\mathrm{Graph}(\mathscr{A})\subseteq P\times_{(\mu,\mathbf{t})}(G_{1}\times H_{1})\times_{(\mu,\mathbf{s})}P$. The source and target maps are the respective projections onto the second and first $P$-factors; thus, we may view $\mathscr{K}$ as a subgroupoid of $\mathscr{G}\times\mathscr{H}^{\mathrm{op}}\times\mathrm{Pair}(P)$. The condition that the tensor fields $J$ and $(N_{1}\times N_{2})\times J$ intertwine the action (Equation (7)) implies that $N\coloneqq(N_{1}\times N_{2})\times(J\times J)$ is a morphism from the tangent Lie groupoid $T\mathscr{K}$ to itself. Thus, $N\in\Omega^{1}(\mathscr{K},T\mathscr{K})$ is a multiplicative $(1,1)$-tensor field on $\mathscr{K}$ covering $J$.
The Lie algebroid of $\mathscr{K}=\left(\mathrm{Graph}(\mathscr{A})\rightrightarrows P\right)$ is the graph of the infinitesimal action $\mathscr{a}:\Gamma\left(A_{\mathscr{G}}\times A_{\mathscr{H}}^{\mathrm{op}}\right)\to\mathfrak{X}(P)$; that is, $A_{\mathscr{K}}=\mathrm{Graph}(\mathscr{a})$. The Lie algebroid $A_{\mathscr{K}}$ is a Lie subalgebroid of $(A_{\mathscr{G}}\times A_{\mathscr{H}}^{\mathrm{op}})\times_{(\rho,\mathrm{d}\mu)}TP\to P$, which integrates to $\mathscr{G}\times\mathscr{H}^{\mathrm{op}}\times\mathrm{Pair}(P)$. Applying the Lie functor to the morphism $N=(N_{1}\times N_{2})\times(J\times J)$ and using Lemma 3.6 to identify $J^{\mathrm{tg}}$ as the infinitesimal counterpart of $J\times J$, we obtain the Lie algebroid morphism $R_{N}=R_{1}\times R_{2}\times J^{\mathrm{tg}}:TA_{\mathscr{K}}\to TA_{\mathscr{K}}$, where $R_{1}$ and $R_{2}$ are the infinitesimal counterparts of $K_{1}$ and $K_{2}$, respectively. The condition that $R_{N}$ preserves $\mathrm{Graph}(\mathscr{a})$ is equivalent to
$\mathrm{d}\mathscr{a}\circ\mu^{*}\left(R_{1}\times R_{2}\right)=J^{\mathrm{tg}}\circ\mathrm{d}\mathscr{a}.$
This equation is one of the requirements of Definition 3.13 (cf. Remark 3.14). It remains to observe that the action $\mathscr{a}$ decomposes as $\mathscr{a}=\mathscr{a}_{1}+\mathscr{a}_{2}$, where $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ are the differentiations of $\mathscr{A}_{1}$ and $\mathscr{A}_{2}$, respectively. By Lemma 3.11, $\mathscr{a}_{1}$ and $\mathscr{a}_{2}$ constitute an infinitesimal Morita $\left(A_{\mathscr{G}},A_{\mathscr{H}}\right)$-bibundle $P$. This completes the proof for multiplicative $(1,1)$-tensor fields.
Regarding the Nijenhuis property, the vanishing of the torsion for the multiplicative tensor fields implies that the IM tensor fields are also Nijenhuis (see [undefg, Section 6.2]). ∎
@@
Let $\mathscr{G}$ and $\mathscr{H}$ be source-simply connected Lie groupoids with Lie algebroids $A_{\mathscr{G}}$ and $A_{\mathscr{H}}$ respectively, and let $P$ be an infinitesimal Morita $(A_{\mathscr{G}},A_{\mathscr{H}})$-bibundle. Let $R_{1}\in\Omega^{1}(A_{\mathscr{G}},TA_{\mathscr{G}})$ and $R_{2}\in\Omega^{1}(A_{\mathscr{H}},TA_{\mathscr{H}})$ be IM tensor fields that are Morita equivalent via $(P,J)$ for some $J\in\Omega^{1}(P,TP)$. Then the multiplicative $(1,1)$-tensor fields $N_{1}$ and $N_{2}$ that integrate $R_{1}$ and $R_{2}$ are Morita equivalent via the associated Morita $(\mathscr{G},\mathscr{H})$-bibundle $(P,J)$. Furthermore, if the infinitesimal Morita equivalence is Nijenhuis, then the global Morita equivalence is also Nijenhuis.
@@
Let $\mathscr{a}_{1}\colon\mu_{1}^{*}A_{\mathscr{G}}\to TP$ and $\mathscr{a}_{2}\colon\mu_{2}^{*}A_{\mathscr{H}}\to TP$ denote the left and right infinitesimal actions of the respective Lie algebroids. The total infinitesimal action is given by $\mathscr{a}(\xi,\zeta)\coloneqq\mathscr{a}_{1}(\xi)+\mathscr{a}_{2}(\zeta)$. Letting $\mu\coloneqq(\mu_{1},\mu_{2})$, Remark 3.14 implies that the tensor fields are Morita equivalent if and only if the image of $T\mathrm{Graph}(\mathscr{a})$ under the map $R\coloneqq\mu^{*}(R_{1}\times R_{2})\times J^{\mathrm{tg}}$ is contained in $T\mathrm{Graph}(\mathscr{a})$. Because each factor of $R$ is a Lie algebroid morphism, $R$ naturally defines a Lie algebroid endomorphism of $T\mathrm{Graph}(\mathscr{a})$.
Let $\mathscr{A}\colon G_{1}\times_{G_{0}}P\times_{H_{0}}H_{1}\to P$ denote the combined global action, defined by $\mathscr{A}(g,p,h)=g\cdot p\cdot h$, which integrates $\mathscr{a}$. The source-simply connected Lie groupoid integrating $\mathrm{Graph}(\mathscr{a})$ is the graph of $\mathscr{A}$. We denote this Lie groupoid by $\mathscr{K}\coloneqq(\mathrm{Graph}(\mathscr{A})\rightrightarrows P)$, which forms a Lie subgroupoid of $\mathscr{G}\times\mathscr{H}^{\mathrm{op}}\times\mathrm{Pair}(P)$. Applying Lie’s second theorem, we integrate the Lie algebroid morphism $R$ to a Lie groupoid morphism $N\colon T\mathscr{K}\to T\mathscr{K}$. Thus, NN constitutes a multiplicative $(1,1)$-tensor field on $\mathscr{K}$.
Let $N_{1}\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ and $N_{2}\in\Omega^{1}(\mathscr{H},T\mathscr{H})$ be the respective integrations of the IM $(1,1)$-tensor fields $R_{1}$ and $R_{2}$. Since Lemma 3.6 guarantees that $J^{\mathrm{tg}}$ integrates to $J\times J$, the uniqueness of integration implies that NN coincides with the restriction of $(N_{1}\times N_{2})\times(J\times J)$ to $\mathrm{Graph}(\mathscr{A})$.
The condition that $(N_{1}\times N_{2})\times(J\times J)$ preserves $\mathrm{Graph}(\mathscr{A})$ is equivalent to the Morita $(\mathscr{G},\mathscr{H})$-bibundle $(P,J)$ establishing a Morita equivalence between the multiplicative tensor fields $N_{1}$ and $N_{2}$. Because the vanishing of the Nijenhuis torsion is preserved under integration (see [undefg, Section 6.2]), the result follows. ∎
@@
4. Compatibility with quasi-symplectic groupoids and Dirac structures
In this section, we apply the framework of Morita equivalence for multiplicative $(1,1)$-tensor fields, developed in Section 2, to Lie groupoids equipped with geometric structures compatible with the Nijenhuis condition. Building upon the established notion of Morita equivalence for quasi-symplectic groupoids introduced by Xu [undefam], our primary objective is to extend this definition to quasi-symplectic-Nijenhuis groupoids and prove that it constitutes an equivalence relation (Theorem 4.6).
Subsequently, we investigate the infinitesimal counterpart of this theory as an application of Section 3. Recall that a Morita equivalence of quasi-symplectic groupoids differentiates to a dual pair of Dirac structures satisfying additional conditions [undefq]. We enhance this correspondence by imposing compatibility with infinitesimal Nijenhuis structures, leading to a notion of Morita equivalence for Dirac-Nijenhuis structures (Definition 4.14). Finally, we establish the consistency of this framework via the Lie functor: Theorems 4.17 and 4.18 prove the global-to-infinitesimal correspondence.
4.1. Quasi-symplectic-Nijenhuis groupoids
We begin by recalling that a quasi-symplectic groupoid [undefam, undefi] (called twisted presymplectic groupoid in [undefi]) is a triple $(\mathscr{G},\omega,\eta)$, where $\mathscr{G}=\left(G_{1}\rightrightarrows G_{0}\right)$ is a Lie groupoid such that $\operatorname{dim}G_{1}=2\operatorname{dim}G_{0}$, $\omega\in\Omega^{2}(G_{1})$ is a multiplicative 22-form and $\eta\in\Omega^{3}(G_{0})$ is a closed 33-form satisfying
(40) $\mathrm{d}\omega=\mathbf{s}^{*}\eta-\mathbf{t}^{*}\eta\qquad\text{and}\qquad\operatorname{ker}\omega\cap\operatorname{ker}\mathrm{d}\mathbf{s}\cap\operatorname{ker}\mathrm{d}\mathbf{t}=0.$
Although Morita equivalence for quasi-symplectic groupoids can be formulated in terms of Morita spans (see [undefad, undefm]), we give preference to the original definition [undefam] here. This is because the formulation via spans does not naturally equip the intertwining groupoid with a multiplicative 22-form, a feature which is essential for stating the compatibility with the Nijenhuis tensor field in our framework.
Recall from [undefam] (see also [undef, Appendix A]) that a Morita equivalence of quasi-symplectic groupoids $\left(\mathscr{G},\omega_{1},\eta_{1}\right)$ and $\left(\mathscr{H},\omega_{2},\eta_{2}\right)$ is given by a Morita bibundle $G_{0}\leftarrow P\to H_{0}$ equipped with a 22-form $\varpi\in\Omega^{2}(P)$,
@@ 4.1. Quasi-symplectic-Nijenhuis groupoids
where $\mathscr{A}\colon G_{1}\times_{G_{0}}P\times_{H_{0}}H_{1}\to P$ is the action $(g,p,h)\mapsto g\cdot p\cdot h$. This definition extends the notion of Morita equivalence for symplectic groupoids introduced by Xu [undefao].
@@
Morita equivalence of quasi-symplectic groupoids is indeed an equivalence relation [undefam, Theorem 4.5]. For transitivity, given two principal bibundles $G_{0}\leftarrow(P_{1},\varpi_{1})\to H_{0}$ and $H_{0}\leftarrow(P_{2},\varpi_{2})\to G_{0}^{\prime}$, a 22-form $\varpi\in\Omega^{2}(P)$ is obtained on the quotient $P=P_{1}\diamond P_{2}=\left(P_{1}\times_{H_{0}}P_{2}\right)/\mathscr{H}$, realizing a Morita equivalence $G_{0}\leftarrow(P,\varpi)\to G_{0}^{\prime}$. The 22-form $\varpi$ is defined by the property
(41) $\pi^{*}\varpi=\mathrm{pr}_{1}^{*}\varpi_{1}+\mathrm{pr}_{2}^{*}\varpi_{2},$
@@
To define quasi-symplectic-Nijenhuis groupoids we require a compatibility between 22-forms and $(1,1)$-tensor fields. Consider the following notation. For any $p$-form $\alpha\in\Omega^{p}\left(M\right)$ and vector-valued 11-form $K\in\Omega^{1}(M,TM)$, write $\alpha_{K}$ to denote
@@
A 22-form $\omega\in\Omega^{2}(M)$ and a $(1,1)$-tensor field $K\in\Omega^{1}(M,TM)$ on a manifold $M$ are said to be compatible [undefac] if
(42) $\omega^{\flat}\circ K=K^{*}\circ\omega^{\flat}\qquad\text{and}$
(43) $\mathrm{d}(\omega_{K})=(\mathrm{d}\omega)_{K}.$
@@
A quasi-symplectic-Nijenhuis groupoid is a quadruple $(\mathscr{G},\omega,\eta,N)$, where $(\mathscr{G},\omega,\eta)$ is a quasi-symplectic groupoid and $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ is a multiplicative Nijenhuis tensor field that is compatible with the 22-form $\omega$.
@@
Following the terminology of [undefh], a holomorphic presymplectic groupoid is a holomorphic groupoid $(\mathscr{G},I_{\mathscr{G}})$ equipped with a multiplicative holomorphic closed 22-form $\Omega\in\Omega^{(2,0)}(G_{1})$ that satisfies $\operatorname{ker}\Omega\cap\operatorname{ker}\mathrm{d}\mathbf{s}\cap\operatorname{ker}\mathrm{d}\mathbf{t}=0$. Equivalently, a holomorphic presymplectic groupoid is described as a quasi-symplectic-Nijenhuis groupoid $(\mathscr{G},\omega,0,I_{\mathscr{G}})$, where $\mathrm{d}\omega=0$ and $\Omega=\omega-i\omega_{I_{\mathscr{G}}}$ [undefh, Section 6.3].
Definition $4.$
A Morita equivalence of quasi-symplectic-Nijenhuis groupoids $(\mathscr{G},\omega_{1},\eta_{1},N_{1})$ and $(\mathscr{H},\omega_{2},\eta_{2},N_{2})$ is given by a Morita $(\mathscr{G},\mathscr{H})$-bibundle $P$ equipped with a 22-form $\varpi\in\Omega^{2}(P)$ and a $(1,1)$-tensor field $J\in\Omega^{1}(P,TP)$,
@@
Two holomorphic presymplectic groupoids $(\mathscr{G},\Omega_{1},I_{\mathscr{G}})$ and $(\mathscr{H},\Omega_{2},I_{\mathscr{H}})$ are Morita equivalent if and only if, there is a holomorphic Morita bibundle $G_{0}\xleftarrow{\ \mu_{1}\ }(P,I_{P})\xrightarrow{\ \mu_{2}\ }H_{0}$ (Example 2.10) and a holomorphic closed 22-form $\Omega_{P}=\varpi-i\varpi_{I_{P}}\in\Omega^{(2,0)}(P)$ satisfying $\mathscr{A}^{*}\varpi=\mathrm{pr}_{G_{1}}^{*}\omega_{1}+\mathrm{pr}_{P}^{*}\varpi+\mathrm{pr}_{H_{1}}^{*}\omega_{2}$ and $\operatorname{ker}\varpi\cap\operatorname{ker}\mathrm{d}\mu_{1}\cap\operatorname{ker}\mathrm{d}\mu_{2}=0$, where $\Omega_{1}=\omega_{1}-i\omega_{1\,I_{\mathscr{G}}}$ and $\Omega_{2}=\omega_{2}-i\omega_{2\,I_{\mathscr{H}}}$.